# Low-Complexity OTFS-Based Over-the-Air Computation Design for Time-Varying Channels

### v1

@@ I Introduction
The latest developments in wireless communication and smart device technologies have spurred the widespread adoption of the Internet of Things (IoT), enabling the seamless connection of billions of physical objects to the Internet through ubiquitous sensing and computing capabilities [2]. According to Dell report [3], there will be 41.6 billion IoT devices in 2025, capable of generating 79.4 zettabytes (ZB) of data. Traditional data aggregation methods via orthogonal channels may not be applicable in future IoT networks due to the restricted spectrum resources. Recently, over-the-air computation (AirComp) was proposed as an efficient method to overcome this issue, which leverages the signal superposition property over multiple access channels (MAC) to allow rapid wireless data aggregation [4]. Due to the unique characteristics of AirComp in integrating computation and communication, its application within IoT networks and distributed machine learning holds a huge potential.
@@ II-A OTFS Modulation & Demodulation
where $g_{\mathrm{tx}}(t)$ is the pulse shaping filter at the TX.
@@
where $R_{u}$ is the number of propagation paths from the $u$-th device to FC, $h_{u,i}$, $\tau_{u,i}$, and $\nu_{u,i}$ represent the path gain, delay and Doppler shifts of the $i$-th path from the $u$-th device to FC, respectively, and $\delta(\cdot)$ is the Dirac delta function. We further assume that the delay and Doppler shifts are the integer multiples of $1/(M\Delta f)$ and $1/(NT)$, respectively, where $NT$ and $M\Delta f$ are the total duration and bandwidth of one OTFS frame, and we have $T\Delta f=1$ [20]. Accordingly, we have $\tau_{u,i}=l_{u,i}/(M\Delta f)$ and $\nu_{u,i}=k_{u,i}/(NT)$, where $l_{u,i}\in\mathcal{L}_{u}$ and $k_{u,i}\in\mathcal{K}_{u}$ are the indices of delay and Doppler. Here, $\mathcal{L}_{u}$ and $\mathcal{K}_{u}$ represent the sets of $l_{u,i}$ and $k_{u,i}$ among all paths between the $u$-th device and FC. Without loss of generality, we assume that $l_{u,1}\leq...\leq l_{u,i}\leq...\leq l_{u,R_{u}}$. Subsequently, the received signal at the FC is given by
@@
In this paper, we consider a practical scenario where both $g_{\mathrm{tx}}(t)$ and $g_{\mathrm{rx}}(t)$ adopt rectangular forms, as in [13, 14]. Under this condition, the input-output relation between $x_{u}[l,k]$ and $y[l,k]$ is given by [20, Eq. (30)]
@@ II-B OTFS-based AirComp System
Compared to the conventional input-output relation of AirComp [9, Eq. (5)], (II-B) suffers from additional ISI and ILI, which are caused by the delay and Doppler spread of multiple paths. From (II-B), we observe that only the time-invariant path gain, delay shift, and Doppler shift of the principal path are required for pre-processing at each device.
@@ III-A Problem Formulation
$\mathcal{P}1:\!\!\!\!\!$ $\min_{p_{u}\geq 0,\eta\geq 0}\sum_{u=1}^{U}\!\!\left(\frac{\sqrt{p_{u}}|h_{u,1}|}{\sqrt{\eta}}-1\right)^{2}\!+\!\sum_{u=1}^{U}\sum_{i=2}^{R_{u}}\frac{p_{u}|h_{u,i}|^{2}}{\eta}+\frac{\sigma^{2}}{\eta},$
$\mathrm{s.t.}\;p_{u}\leq P_{\mathrm{s}},\forall u\in\mathcal{U}.$ (14)
@@
Filters with ideal waveforms, i.e., $g_{\mathrm{tx}}(t)$ and $g_{\mathrm{rx}}(t)$, which satisfy the bi-orthogonal property [12, Eq. (4)] at the TX and RX, represent another widely considered scenario in OTFS modulation. According to [20, Eq. (6)], the input-output relation between $x_{u}[l,k]$ and $y[l,k]$ under this condition is given by
@@ III-B Optimization of $p_{u}$ and $\eta$
$\mathcal{P}2:$ $\min_{\eta\geq 0}\sum_{u=1}^{U}\left(\min\left(\frac{\sqrt{P_{\mathrm{s}}}|h_{u,1}|}{\sqrt{\eta}},\frac{|h_{u,1}|^{2}}{\sum_{i=1}^{R_{u}}|h_{u,1}|^{2}}\right)-1\right)^{2}$
@@ IV-A ZP-Assisted OTFS
In particular, $\boldsymbol{K}^{u}_{m,l}$ occupies $\tilde{\boldsymbol{H}}_{u}$ within the rows from $mN+1$ to $(m+1)N$ and columns from $([m+1-l]_{M}-1)N+1$ to $[m+1-l]_{m}N$. According to (27) and Fig. 2, it is evident that the configuration of $\tilde{\boldsymbol{H}}_{u}$ is influenced by both the number of channel paths and the respective delay and Doppler shifts within each path. For different typologies of $\tilde{\boldsymbol{H}}_{u}$ at each device, we need to arrange the sequence of data accordingly to ensure that the corresponding $\boldsymbol{x}_{u}$ can be aggregated. In this section, we consider a scenario that $R_{1}=...=R_{U}=R$, $l_{1,i}=...=l_{U,i}=l_{i}$, and $k_{1,i}=...=k_{U,i}=k_{i}$. We note that ensuring a uniform Doppler shift requires maintaining an identical velocity of all devices relative to the FC, and an equal number of paths can be achieved when the environment between the FC and each device is relatively homogeneous. We acknowledge that if these conditions do not hold, the corresponding signals cannot be properly aggregated and additional interference exists, degrading the effect of proposed SIC-based algorithm. Then, we can rearrange the data mapping on the DD domain according to delays to ensure that the corresponding signals are aggregated and estimate the target function with the existence of additional interference as in Section III. The details of estimation methods design for applying ZP-assisted OTFS in these scenarios will be investigated in the future.
@@ IV-B1 Different $l_{i}$ associated to each path
where $\sqrt{p_{u,m}}$ is the transmit power for $x_{u,\gamma_{m}}$. We denote $\hat{f}_{m}$ as the estimated value of $\sum_{u=1}^{U}x_{u,\gamma_{m}}$ and $\epsilon_{m}$ as the MSE associated with estimating $\sum_{u=1}^{U}x_{u,\gamma_{m}}$. Here, $\hat{f}_{0}=y_{\gamma_{0}}/\sqrt{\eta_{0}}$, where $\eta_{m}$ is the denoising factor for $y_{\gamma_{m}}$. We note that (31) has the same format as that of [9, Eq. (5)], which suggests that the optimization procedures of $p_{u,0}$ and $\eta_{0}$ can follow [9]. Thus, $p_{u,0}^{*}$ and $\eta_{0}^{*}$ are given by [9, Eqs. (24), (25)]
@@
$y_{\gamma_{m_{2}}}\!\!=\!\!\sum_{u=1}^{U}\!\sqrt{p_{u,m_{1}}}|\nu_{m_{2},l_{1}}^{u}|x_{u,\gamma_{m_{1}}}\!\!+\!\!\!\sum_{m\in\mathcal{M}}\!\sum_{u=1}^{U}\!\sqrt{p_{u,m}^{*}}\nu_{m_{2},m_{1}-m+l}^{u}$
×$\times\frac{(\nu_{m_{2},l_{1}}^{u})^{\dagger}x_{u,\gamma_{m}}}{|\nu_{m_{2},l_{1}}^{u}|}+w_{\gamma_{m_{2}}}-\sum_{m\in\mathcal{M}}\zeta_{m_{2},m}\hat{f}_{m},$ (45)
@@ IV-B2 At least two paths share the same $l$
$+\!\!\sum_{u=1}^{U}\!\sum_{\varphi\in\mathcal{G}_{l_{1}}\setminus\{\varphi_{1}\}}\!\!\!\!\!\!\!\!\!\!\frac{p_{u,m_{1}}\!|\boldsymbol{\nu}_{m_{2},l}^{u}(\varphi)|^{2}}{\eta_{m_{2}}}\!+\!\frac{\sigma^{2}}{\eta_{m_{2}}},\mathrm{s.t.}\;p_{u,m_{1}}\!\leq\!P_{\mathrm{s}},\forall u\!\in\mathcal{U}.$ (48)
@@ V-A Problem Formulation
We denote $\boldsymbol{x}_{u}\in\mathbb{C}^{MN\times 1}$ and $\boldsymbol{y}\in\mathbb{C}^{MN\times 1}$ as the column-wise vectorization of $\boldsymbol{X}_{u}$ and $\boldsymbol{Y}$. Then, (7) can be rephrased in a matrix format as $\boldsymbol{y}=\sum_{u=1}^{U}\boldsymbol{H}_{u}\boldsymbol{x}_{u}+\boldsymbol{w}$ [20, Eq. (11)], where $\boldsymbol{H}_{u}=(\boldsymbol{F}_{N}\otimes\boldsymbol{I}_{M})\hat{\boldsymbol{H}}_{u}(\boldsymbol{F}_{N}^{H}\otimes\boldsymbol{I}_{M})$ with $\boldsymbol{F}_{N}\in\mathbb{C}^{N\times N}$ represents an $N$-point FFT, $\boldsymbol{F}_{N}^{H}$ is an $N$-point IFFT, $\tilde{\boldsymbol{w}}\in\mathbb{C}^{MN\times 1}$ is the noise vector, and $\hat{\boldsymbol{H}}_{u}$ is given by $\hat{\boldsymbol{H}}_{u}=\sum_{i=1}^{R_{u}}h_{u,i}\boldsymbol{\Pi}^{l_{u,i}}\boldsymbol{\Delta}^{k_{u,i}}$, where $\boldsymbol{\Pi}\in\mathbb{C}^{MN\times MN}$ is the permutation matrix given by
@@
$\epsilon\!=\!\!\!\sum_{u=1}^{U}\!\mathrm{Tr}\!\left[\!\left(\boldsymbol{V}\!\boldsymbol{H}_{u}\!\boldsymbol{B}_{u}\!-\!\boldsymbol{I}\right)\!\left(\boldsymbol{V}\!\boldsymbol{H}_{u}\!\boldsymbol{B}_{u}\!-\boldsymbol{I}\right)^{H}\right]+\sigma^{2}\mathrm{Tr}\left(\boldsymbol{V}\boldsymbol{V}^{H}\right),$ (54)
where $\mathrm{Tr}(\cdot)$ represents trace and $\boldsymbol{V}^{H}$ represents the Hermitian transpose of $\boldsymbol{V}$. Our objective is to minimize (54) by jointly optimizing $\boldsymbol{V}$ and $\boldsymbol{B}_{u}$. We assume that $P_{\mathrm{t}}$ is the total average power budget for transmitting $MN$ data at each device, and we have $P_{\mathrm{t}}=MNP_{\mathrm{s}}$. Then, we formulate the optimization problem, referred to as $\mathcal{P}3$, as follows
@@ V-B Optimization of $\boldsymbol{B}_{u}$ and $\boldsymbol{V}$
where the inverse of $\boldsymbol{\Sigma}_{u}+\lambda_{u}^{*}\boldsymbol{I}$ can be directly obtained by taking the reciprocal of the diagonal element.
@@ VI Numerical Results
In this section, we evaluate the MSE performance and computational complexity of S1, S2, and S3. For S1, the MSE is computed by applying the optimal transmit power and the optimal denoising factor derived in Theorem 1. For S2, the MSE is computed based on the SIC-based algorithm, where the optimal transmit power, denoising factor, and the weight for interference cancellation are derived in Theorem 2. For S3, the MSE is computed based on the optimized precoding matrix and receiving filter matrix in Algorithm 1. Moreover, the evaluation focuses on the average MSE over a large number of channel realizations, which we set 1000 in the simulation. The path gains are randomly generated based on a uniform power delay profile, and the delay and Doppler indices are randomly generated within the range of $[0,l_{\mathrm{max}}]$ and $[-k_{\mathrm{max}},k_{\mathrm{max}}]$. Specifically, we set $M=32$, $N=16$, $U=20$, $\Delta f=1.5\;\mathrm{kHz}$, carrier frequency $f_{\mathrm{c}}=4\;\mathrm{GHz}$, $l_{\mathrm{max}}=10$, and $k_{\mathrm{max}}=5$ [23], unless otherwise stated.
@@
  • The first benchmark is proposed in [17], where the authors set the precoding matrix as $\boldsymbol{B}_{u}=(\boldsymbol{H}_{u}^{H}\boldsymbol{H}_{u}+\sigma^{2}\boldsymbol{I})^{-1}\boldsymbol{H}_{u}^{H}$, and $\hat{\boldsymbol{f}}$ is given by $\hat{\boldsymbol{f}}=\sum_{u=1}^{U}(\boldsymbol{H}_{u}\boldsymbol{B}_{u}\boldsymbol{x}_{u}+\boldsymbol{w}_{u}/\phi_{u})$, where $\phi_{u}=\sqrt{P_{\mathrm{t}}/\mathrm{Tr}(\boldsymbol{B}_{u}\boldsymbol{B}_{u}^{H})}$ is the power normalization factor.
  • The second benchmark [18] is proposed based on [17], where the authors include power control at each device. In particular, the authors replace $P_{\mathrm{t}}$ in $\phi_{u}$ with a transmit power $p_{u}$, $p_{u}\leq P_{\mathrm{t}}$, and optimize $p_{u}$ to minimize the MSE.
  • The third benchmark is a simplified version of S3, where we only consider a precoding matrix at each device. By setting $\boldsymbol{V}^{*}=\boldsymbol{I}$ in (61), we obtain $\boldsymbol{B}_{u}^{*}$ as $\boldsymbol{B}_{u}^{*}=(\boldsymbol{H}_{u}^{H}\boldsymbol{H}_{u}+\lambda_{u}^{*}\boldsymbol{I})^{-1}\boldsymbol{H}_{u}^{H}$.
  • The fourth benchmark only considers a receiving filter matrix at FC. By setting $\boldsymbol{B}_{u}^{*}=\boldsymbol{I}$ in (62), we obtain $\boldsymbol{V}^{*}$ as $\boldsymbol{V}^{*}=\left(\sum_{u=1}^{U}\boldsymbol{H}_{u}^{H}\right)\left(\sum_{u=1}^{U}\boldsymbol{H}_{u}\boldsymbol{H}_{u}^{H}+\sigma^{2}\boldsymbol{I}\right)^{-1}$.

### v2

@@ I Introduction
The latest developments in wireless communication and smart device technologies have spurred the widespread adoption of the Internet of Things (IoT), enabling the seamless connection of billions of physical objects to the Internet through ubiquitous sensing and computing capabilities [2]. According to Dell report [3], there will be 41.6 billion IoT devices in 2025, capable of generating 79.4 zettabytes (ZB) of data. Traditional data aggregation methods via orthogonal channels may not be applicable in future IoT networks due to the restricted spectrum resources. Recently, over-the-air computation (AirComp) was proposed as an efficient method to overcome this issue, which leverages the signal superposition property over multiple access channels (MAC) to allow rapid wireless data aggregation [4]. Due to the unique characteristics of AirComp in integrating computation and communication, its application in IoT networks, distributed machine learning, and AI-assisted autonomous driving systems holds significant potential [5].
@@ II-A OTFS Modulation & Demodulation
where $g_{\mathrm{tx}}(t)$ is the pulse shaping filter at the TX, $T$ represents the duration of each time slot, and $\Delta f$ is the subcarrier spacing. To ensure TF symbols are orthogonal to each other, we have $T\Delta f=1$.
@@
where $R_{u}$ is the number of propagation paths from the $u$-th device to FC, $h_{u,i}$, $\tau_{u,i}$, and $\nu_{u,i}$ represent the path gain, delay and Doppler shifts of the $i$-th path from the $u$-th device to FC, respectively, and $\delta(\cdot)$ is the Dirac delta function. We further assume that the delay and Doppler shifts are the integer multiples of $1/(M\Delta f)$ and $1/(NT)$, respectively, where $NT$ and $M\Delta f$ are the total duration and bandwidth of one OTFS frame [21]. We clarify that the fractional Doppler shift can be addressed by adding virtual integer taps in the DD domain channel [22] or by applying TF domain windows [23]. In practice, the non-fractional case can be achieved by using sufficiently large $M$ and $N$ [14]. Accordingly, we have $\tau_{u,i}=l_{u,i}/(M\Delta f)$ and $\nu_{u,i}=k_{u,i}/(NT)$, where $l_{u,i}\in\mathcal{L}_{u}$ and $k_{u,i}\in\mathcal{K}_{u}$ are the indices of delay and Doppler. Here, $\mathcal{L}_{u}$ and $\mathcal{K}_{u}$ represent the sets of $l_{u,i}$ and $k_{u,i}$ among all paths between the $u$-th device and FC. Given the velocity of the $u$-th device as $v_{u}$, $k_{u,i}$ can be computed as
$k_{u,i}=\frac{v_{u}f_{\mathrm{c}}\cos(\rho_{u,i})N}{c\Delta f},$ (4)
where $f_{\mathrm{c}}$ is the carrier frequency, $\rho_{u,i}$, is the angle between the wave propagation direction of the $i$-th path and the motion direction of the $u$-th device, and $c$ is the speed of light. Without loss of generality, we assume that $l_{u,1}\leq...\leq l_{u,i}\leq...\leq l_{u,R_{u}}$. Subsequently, the received signal at the FC is given by
@@
In this paper, we consider a practical scenario where both $g_{\mathrm{tx}}(t)$ and $g_{\mathrm{rx}}(t)$ adopt rectangular forms, as in [14, 15]. Under this condition, the input-output relation between $x_{u}[l,k]$ and $y[l,k]$ is given by [21, Eq. (30)]
@@ II-B OTFS-based AirComp System
Compared to the conventional input-output relation of AirComp [10, Eq. (5)], (II-B) suffers from additional ISI and ILI, which are caused by the delay and Doppler spread of multiple paths. From (II-B), we observe that only the time-invariant path gain, delay shift, and Doppler shift of the principal path are required for pre-processing at each device.
@@
We note that calculating the average function is essential in many applications. For instance, in intelligent transportation systems, vehicles share information like speed, position, and environmental conditions to compute the average speed or density of vehicles on a particular road segment in real-time. This can improve traffic light control, dynamic tolling, and congestion avoidance by reducing the need for centralized data processing. Similarly, in federated learning, mobile devices, e.g., smartphones, vehicles, or wearables, train local machine learning models and aggregate updates at the server. The server computes the average of the model gradients to update the global model.
@@ III-A Problem Formulation
$\mathcal{P}1:$ $\min_{p_{u}\geq 0,\eta\geq 0}\sum_{u=1}^{U}\left(\frac{\sqrt{p_{u}}|h_{u,1}|}{\sqrt{\eta}}-1\right)^{2}+\sum_{u=1}^{U}\sum_{i=2}^{R_{u}}\frac{p_{u}|h_{u,i}|^{2}}{\eta}$
$+\frac{\sigma^{2}}{\eta},\;\;\mathrm{s.t.}\;p_{u}\leq P_{\mathrm{s}},\forall u\in\mathcal{U}.$ (15)
@@
Filters with ideal waveforms, i.e., $g_{\mathrm{tx}}(t)$ and $g_{\mathrm{rx}}(t)$, which satisfy the bi-orthogonal property [13, Eq. (4)] at the TX and RX, represent another widely considered scenario in OTFS modulation. According to [21, Eq. (6)], the input-output relation between $x_{u}[l,k]$ and $y[l,k]$ under this condition is given by
@@ III-B Optimization of $p_{u}$ and $\eta$
$\mathcal{P}2:$ $\min_{\eta\geq 0}\sum_{u=1}^{U}\left(\min\left(\frac{\sqrt{P_{\mathrm{s}}}|h_{u,1}|}{\sqrt{\eta}},\frac{|h_{u,1}|^{2}}{\sum_{i=1}^{R_{u}}|h_{u,i}|^{2}}\right)-1\right)^{2}$
@@
From Theorem 1, it is evident that path gains are essential at both the FC and each device. In practical scenarios, the FC can estimate channel parameters based on the signals transmitted by the devices. This can be achieved by arranging pilot symbols on the DD grid, alongside guard and data symbols. The FC can then utilize methods such as the threshold approach [24] or off-grid sparse Bayesian learning [25] to estimate the channel information. Once the channel information is acquired, the FC transmits it back to the devices, enabling them to perform the pre-processing.
@@ IV-A ZP-Assisted OTFS
In particular, $\boldsymbol{K}^{u}_{m,l}$ occupies $\tilde{\boldsymbol{H}}_{u}$ within the rows from $mN+1$ to $(m+1)N$ and columns from $([m+1-l]_{M}-1)N+1$ to $[m+1-l]_{m}N$. According to (28) and Fig. 2, it is evident that the configuration of $\tilde{\boldsymbol{H}}_{u}$ is influenced by Doppler shifts at each path. For different typologies of $\tilde{\boldsymbol{H}}_{u}$ at each device, we need to arrange the sequence of data accordingly to ensure that the corresponding $\boldsymbol{x}_{u}$ can be aggregated. In this section, we explore a scenario where all devices experience the same Doppler shift, allowing efficient SIC. This situation can occur when devices move together as a group, ensuring they travel in the same environment at the same speed. A practical example of this is multiple devices, such as smartphones, being on the same vehicle when transmitting data to the FC. Another relevant case is in satellite communications, where a satellite serves as the FC, receiving signals from ground-based devices. As satellites typically move at very high velocities, e.g., $7-8\;\mathrm{km}/\mathrm{s}$, it dominates the relative velocity between the satellite and the devices on the ground, given that the speed of ground vehicles is usually around $20-100\;\mathrm{m/s}$. Therefore, we can regard a same Doppler shift for all devices if they are moving in a similar direction relative to the satellite’s motion, e.g., both moving towards or away from the satellite [26].
@@ IV-B1 Different $l_{i}$ associated to each path
where $\sqrt{p_{u,m}}$ is the transmit power for $x_{u,\gamma_{m}}$. We denote $\hat{f}_{m}$ as the estimated value of $\sum_{u=1}^{U}x_{u,\gamma_{m}}$ and $\epsilon_{m}$ as the MSE associated with estimating $\sum_{u=1}^{U}x_{u,\gamma_{m}}$. Here, $\hat{f}_{0}=y_{\gamma_{0}}/\sqrt{\eta_{0}}$, where $\eta_{m}$ is the denoising factor for $y_{\gamma_{m}}$. We note that (32) has the same format as that of [10, Eq. (5)], which suggests that the optimization procedures of $p_{u,0}$ and $\eta_{0}$ can follow [10]. Thus, $p_{u,0}^{*}$ and $\eta_{0}^{*}$ are given by [10, Eqs. (24), (25)]
@@
$y_{\gamma_{m_{2}}}=\sum_{u=1}^{U}\sqrt{p_{u,m_{1}}}|\nu_{m_{2},l_{1}}^{u}|x_{u,\gamma_{m_{1}}}$
$+\sum_{m\in\mathcal{M}}\sum_{u=1}^{U}\sqrt{p_{u,m}^{*}}\nu_{m_{2},m_{1}-m+l}^{u}\frac{(\nu_{m_{2},l_{1}}^{u})^{\dagger}x_{u,\gamma_{m}}}{|\nu_{m_{2},l_{1}}^{u}|}+w_{\gamma_{m_{2}}}$
$-\sum_{m\in\mathcal{M}}\zeta_{m_{2},m}\hat{f}_{m},$ (46)
@@
Based on Theorem 2, it is evident that in addition to requiring the channel gains at both the FC and the devices to determine the optimal transmit power, denoising factor, and interference cancellation weights, the FC also needs the optimal transmit power used for transmitting the previous symbols. This necessitates that each device communicates its optimal transmit power to the FC. Furthermore, once the FC calculates the optimal denoising factor, it relays this information to all devices, enabling them to compute their respective optimal transmit powers.
@@ IV-B2 At least two paths share the same $l$
$+\sum_{u=1}^{U}\sum_{\varphi\in\mathcal{G}_{l_{1}}\setminus\{\varphi_{1}\}}\frac{p_{u,m_{1}}|\boldsymbol{\nu}_{m_{2},l}^{u}(\varphi)|^{2}}{\eta_{m_{2}}}+\frac{\sigma^{2}}{\eta_{m_{2}}},$
$\mathrm{s.t.}\;p_{u,m_{1}}\leq P_{\mathrm{s}},\forall u\in\mathcal{U}.$ (49)
@@ V-A Problem Formulation
We denote $\boldsymbol{x}_{u}\in\mathbb{C}^{MN\times 1}$ and $\boldsymbol{y}\in\mathbb{C}^{MN\times 1}$ as the column-wise vectorization of $\boldsymbol{X}_{u}$ and $\boldsymbol{Y}$. Then, (8) can be rephrased in a matrix format as $\boldsymbol{y}=\sum_{u=1}^{U}\boldsymbol{H}_{u}\boldsymbol{x}_{u}+\boldsymbol{w}$ [21, Eq. (11)], where $\boldsymbol{H}_{u}=(\boldsymbol{F}_{N}\otimes\boldsymbol{I}_{M})\hat{\boldsymbol{H}}_{u}(\boldsymbol{F}_{N}^{H}\otimes\boldsymbol{I}_{M})$ with $\boldsymbol{F}_{N}\in\mathbb{C}^{N\times N}$ represents an $N$-point FFT, $\boldsymbol{F}_{N}^{H}$ is an $N$-point IFFT, $\tilde{\boldsymbol{w}}\in\mathbb{C}^{MN\times 1}$ is the noise vector, and $\hat{\boldsymbol{H}}_{u}$ is given by $\hat{\boldsymbol{H}}_{u}=\sum_{i=1}^{R_{u}}h_{u,i}\boldsymbol{\Pi}^{l_{u,i}}\boldsymbol{\Delta}^{k_{u,i}}$, where $\boldsymbol{\Pi}\in\mathbb{C}^{MN\times MN}$ is the permutation matrix given by
@@
ϵ=italic-$\epsilon=$ $\frac{1}{U^{2}}\left[\sum_{u=1}^{U}\mathrm{Tr}\left[\left(\boldsymbol{V}\boldsymbol{H}_{u}\boldsymbol{B}_{u}-\boldsymbol{I}\right)\left(\boldsymbol{V}\boldsymbol{H}_{u}\boldsymbol{B}_{u}-\boldsymbol{I}\right)^{H}\right]\right.$
$\left.+\sigma^{2}\mathrm{Tr}\left(\boldsymbol{V}\boldsymbol{V}^{H}\right)\right],$ (55)
where $\mathrm{Tr}(\cdot)$ represents trace and $\boldsymbol{V}^{H}$ represents the Hermitian transpose of $\boldsymbol{V}$. Our objective is to minimize (V-A) by jointly optimizing $\boldsymbol{V}$ and $\boldsymbol{B}_{u}$. We assume that $P_{\mathrm{t}}$ is the total average power budget for transmitting $MN$ data at each device, and we have $P_{\mathrm{t}}=MNP_{\mathrm{s}}$. Then, we formulate the optimization problem, referred to as $\mathcal{P}3$, as follows
@@ V-B Optimization of $\boldsymbol{B}_{u}$ and $\boldsymbol{V}$
where the inverse of $\boldsymbol{\Sigma}_{u}+\lambda_{u}^{*}\boldsymbol{I}$ can be directly obtained by taking the reciprocal of the diagonal element. Additionally, the computational complexity of each iteration in Algorithm 1 is $\mathcal{O}(U(M^{3}N^{3}+M^{2}N^{2}))$.
@@ VI Numerical Results
In this section, we evaluate the MSE performance and computational complexity of S1, S2, and S3. For S1, the MSE is computed by applying the optimal transmit power and the optimal denoising factor derived in Theorem 1. For S2, the MSE is computed based on the SIC-based algorithm, where the optimal transmit power, denoising factor, and the weight for interference cancellation are derived in Theorem 2. For S3, the MSE is computed based on the optimized precoding matrix and receiving filter matrix in Algorithm 1. Moreover, the evaluation focuses on the average MSE over a large number of channel realizations, which we set 1000 in the simulation. The path gains are randomly generated based on a uniform power delay profile, and the delay and Doppler indices are randomly generated within the range of $[0,l_{\mathrm{max}}]$ and $[-k_{\mathrm{max}},k_{\mathrm{max}}]$. Specifically, we set $M=32$, $N=16$, $U=20$, $\Delta f=1.5\;\mathrm{kHz}$, $f_{\mathrm{c}}=4\;\mathrm{GHz}$, $l_{\mathrm{max}}=10$, and $k_{\mathrm{max}}=5$ [29], unless otherwise stated.
@@
  • The first benchmark is proposed in [18], where the authors set the precoding matrix as $\boldsymbol{B}_{u}=(\boldsymbol{H}_{u}^{H}\boldsymbol{H}_{u}+\sigma^{2}\boldsymbol{I})^{-1}\boldsymbol{H}_{u}^{H}$, and $\hat{\boldsymbol{f}}$ is given by $\hat{\boldsymbol{f}}=\sum_{u=1}^{U}(\boldsymbol{H}_{u}\boldsymbol{B}_{u}\boldsymbol{x}_{u}+\boldsymbol{w}_{u}/\phi_{u})$, where $\phi_{u}=\sqrt{P_{\mathrm{t}}/\mathrm{Tr}(\boldsymbol{B}_{u}\boldsymbol{B}_{u}^{H})}$ is the power normalization factor. Then, MSE is calculated as
    ϵ=italic-$\epsilon=$ $\frac{1}{U^{2}}\left[\sum_{u=1}^{U}\mathrm{Tr}\left[\left(\boldsymbol{H}_{u}\boldsymbol{B}_{u}-\boldsymbol{I}\right)\left(\boldsymbol{H}_{u}\boldsymbol{B}_{u}-\boldsymbol{I}\right)^{H}\right]\right.$
    $\left.+\frac{\sigma^{2}\mathrm{Tr}(\boldsymbol{B}_{u}\boldsymbol{B}_{u}^{H})}{P_{t}}\right].$ (66)
  • The second benchmark [19] is proposed based on [18], where the authors include power control at each device. In particular, MSE is calculated by replacing $P_{\mathrm{t}}$ with $p_{u}$, $p_{u}\leq P_{\mathrm{t}}$, in (VI), and optimal $p_{u}$ is obtained by exhavsive search to achieve the minimum MSE.
  • The third benchmark is a simplified version of S3, where we only consider a precoding matrix at each device. By setting $\boldsymbol{V}^{*}=\boldsymbol{I}$ in (62), we obtain $\boldsymbol{B}_{u}^{*}$ as $\boldsymbol{B}_{u}^{*}=(\boldsymbol{H}_{u}^{H}\boldsymbol{H}_{u}+\lambda_{u}^{*}\boldsymbol{I})^{-1}\boldsymbol{H}_{u}^{H}$. MSE is obtained by setting $\boldsymbol{V}$ as $\boldsymbol{I}$ in (V-A).
  • The fourth benchmark only considers a receiving filter matrix at FC. By setting $\boldsymbol{B}_{u}^{*}=\boldsymbol{I}$ in (63), we obtain $\boldsymbol{V}^{*}$ as $\boldsymbol{V}^{*}=\left(\sum_{u=1}^{U}\boldsymbol{H}_{u}^{H}\right)\left(\sum_{u=1}^{U}\boldsymbol{H}_{u}\boldsymbol{H}_{u}^{H}+\sigma^{2}\boldsymbol{I}\right)^{-1}$. MSE is obtained by setting $\boldsymbol{B}_{u}$ as $\boldsymbol{I}$ in (V-A).